bracefaceniya1:

Sean says that to add a number to –100 and still have –100 is to add zero. Candice says that she can add two numbers to –100 and still have –100. Who is correct and why?

Isolation:

Have you thought yet that they both may be correct?

TETSXPREME:

@Bracefaceniya1 Welcome to Questioncove. Do Have any thoughts about this?

jhonyy9:

1. Sean says -100 = -100 +0 2. Candice says -100 = -100 +5 + (-5) any idea now ?